While completing your research on the speciation of clover in a county, you isolate an acid from clover leaves. You discover that it contains only the elements C, H, and O. The apparatus is used for analysis of 0.513 g of the compound. Before the experiment, the MgSO4 has a mass of 56.784 g. After the experiment, the MgSO4 has become hydrated to a mass of 56.887 g. Similarly, the mass of NaOH has increased to 34.302 g from 33.801 g. Another experiment has shown that the molar mass of the acid is about 90 g/mol. What is the molecular formula of the acid?
1
Expert's answer
2018-01-25T12:25:08-0500
Since MgSO4 is a dehydrating agent, the mass difference represents mass of water formed due to the combustion reaction,
m(H2O) = 56.887-56.784 = 0.103 g.
The mass of NaOH was increased due to the absorption of CO2 formed in the combustion reaction:
m(CO2) = 34.302-33.801 = 0.501 g.
Having these masses, it is possible to determine the grams of carbon in 0.501 g CO2 and the grams of hydrogen in 0.103 H2O.
carbon: 0.501 g x (12.011 g / 44.0098 g) = 0.1367 g,
hydrogen: 0.103 g x (2.0158 g / 18.0152 g) = 0.01153 g.
The mass of oxygen in the sample can be determined by subtraction:
0.513 g-0.1367 g-0.01153 g =0.365 g.
Converting grams of C, H and O to their respective amount of moles:
carbon: 0.1367 g / 12.011 g/mol = 0.0114 mol,
hydrogen: 0.01153 g / 1.0079 g/mol = 0.0114 mol,
oxygen: 0.365 g / 15.9994 g/mol = 0.0228 mol.
Dividing each molar amount by the lowest value it is possible to modify the molar amounts into small, whole numbers:
0.0114: 0.0114: 0.0228 =1:1:2,
therefore, the empirical formula is CHO2. The empirical formula mass is 12+1+32 = 45 g/mol , which is two times less than the actual molar mass 90 g/mol, the scaling factor is 2, and the molecular formula must be C2H2O4.
Answer: C2H2O4.